# From Nodes to Narratives: Explaining Graph Neural Networks with LLMs and Graph Context

### v1

@@ 5.4 Qualitative Evaluation
Results on Cora. We follow the qualitative evaluation framework provided in Cedro and Martens (2025), and refer to Appendix F for a full description of the evaluation dimensions. An expert reviewed the explanations from GNNexplainer and from our method over 10 node classification explanation tasks, and the evaluation framework is developed in Appendix G. The results are presented in Table 5.3. We observe that our method appeared more trustworthy and convincing to our expert, which may be explained by the semantic analysis abilities of the LLM of our graph explaining method. Concurrently, the explanations were also less insightful, as this may have limited the LLM in its exploration of topics and application not directly limited to the subject of the research paper.
@@
Visualization on Amazon dataset. Figure 2 illustrates a representative example of Logic output on the Amazon dataset to emphasize the narrative quality and interpretability of the generated explanation. In this instance, the target (bold) node corresponds to a product (ID 7), predicted to belong to the “Sports & Outdoors” category. Logic analyzes the local neighborhood of the target node and translates the latent GNN-based representations into natural language explanations by leveraging the LLM’s language generation capabilities, interleaved with soft GNN prompts.
Logic produces node-wise rationales from the computation tree of the target node, explicitly stating for each neighbor a brief semantic summary along with a binary support judgment (YES/NO) regarding its influence on the target’s classification.
Cora WikiCS Liar Amazon
Fidelity Size Fidelity Size Fidelity Size Fidelity Size
LLM 93.3% 20.4 91.2% 9.8 95.0% 11.1 86.5% 5.2
LLM+Pr 92.2 % 16.0 80.1% 3.7 95.0% 10.1 86.5% 1.7
LLM+Pr+Po 94.6% 17.4 89.0% 8.9 94.8% 10.3 87.0% 4.9
5.5 Ablation Study
To assess the effectiveness of the different components of our approach, we measure the performance of our method with and without the projector and the post-processing. The results are in Table 3. We observe that both fidelity and explanation size drop across all datasets when we replace the words in the prompt with the projected GNN embeddings. This might be due to the LLM having to extract knowledge from more abstract representations, making it more selective in its reasoning. We also observe that adding the post-processing step increases fidelity, at a reduced cost in size compared to the method without a projector, on the WikiCS dataset. From this analysis, we suggest that our method offers a good compromise: achieving high fidelity (slightly lower than directly prompting the LLM with node-associated text) while maintaining a compact explanation size (slightly larger than projected GNN embeddings without post-processing).
@@ 6 Conclusions
In this work, we propose Logic, a lightweight, post-hoc explanation framework that uses LLMs to generate faithful and human-interpretable explanations for GNN predictions on text-attributed graphs. By projecting GNN embeddings into the LLM space, Logic enables context-aware, natural language rationales without relying on traditional GNN explainer modules that often need training. Our approach produces concise explanation subgraphs with high fidelity and semantic clarity, outperforming prior methods. Logic demonstrates that LLMs can serve as effective interpreters for GNNs and offers a scalable and intuitive path toward explainability in graph-based applications.
@@ 8 Limitations
  • Our method projects GNN embeddings into the LLM embedding space and interleaves them with text to form hybrid prompts. However, batching multiple nodes leads to degraded performance and hallucinations, likely due to interference between soft prompts.
  • While Logic achieves high empirical fidelity, it does not provide formal guarantees of faithfulness to the GNN’s internal decision logic, as the explanation relies on the LLM’s interpretation of projected embeddings rather than directly tracing GNN computation paths.
  • Although our method incorporated a projector to align the embeddings of the GNN with the LLM space of token embeddings, the objectives we introduce for training the projector are not directly tailored for the task of explanation. More optimal objectives that adhere to the explanation task could bring substantial benefits.
@@ Appendix A Additional Related Work
Here we provide additional related works that are on the integrated methods of LLMs and GNNs. The integration of LLMs with GNNs aims to leverage the advanced reasoning abilities of LLMs and the extensive background knowledge they acquire, alongside the GNNs’ proven capacity to enhance the exploitation of graph structures via message-passing mechanisms Kipf and Welling (2017b); Hamilton et al. (2017). The approaches for integrating LLMs and GNNs primarily vary in the sequence of component application Ren et al. (2024); Chen et al. (2024b).
@@
LLMs as prefix and postfix. First, there are methods that utilize LLMs at the prefix modules of their frameworks for task and feature representation Liu et al. (2024a); Fang et al. (2024); He et al. (2024a); Yang et al. (2025). This architecture provides greater flexibility, as various tasks can be specified in natural language with text-based graph representations, and LLM modules can remain fixed. On the other hand, the second class of methods uses GNNs as the prefix Wang et al. (2023); Chen et al. (2024a); Hu et al. (2024); Tang et al. (2024), which is better suited for capturing a graph’s topology and its intricate relationships and uses LLMs for prediction. This approach also offers flexibility, as the LLM can output embeddings, prediction scores, or textual sequences to address various downstream problems specifically for generation tasks Wang et al. (2023); Chen et al. (2024a); Hu et al. (2024); Tang et al. (2024).
@@
Independent LLMs and joint LLM-GNNs. Additionally, there is a body of work that employs LLMs independently to tackle graph-related tasks Cheng et al. (2024); Fatemi et al. (2024); Li et al. (2024). Approaches following this paradigm benefit from the generalization capabilities of LLMs at both input and output levels, and can potentially avoid training learnable parameters by fixing the LLM component and relying on input prompts. Finally, there is a group of works which employ Gnns and LLMs jointly through soft prompting, which necessitates alignments between their embeddings while enhancing LLMs reasoning by the Gnn captured graph structural features Liu et al. (2024b); Wang et al. (2024); Tian et al. (2023).
@@
Explainers for GNNs. Several algorithms have been proposed in the literature to enhance the interpretability of GNNs. The majority of existing explainers concentrate on instance-level explanations. Instance-level (or local) explainers operate on individual graphs, identifying influential components—such as subgraphs—that most strongly impact the model’s prediction Ying et al. (2019b); Luo et al. (2020); Shan et al. (2021); Huang et al. (2022); Yuan et al. (2022); Lucic et al. (2022b); Tan et al. (2022b); Lin et al. (2021); Bajaj et al. (2021); Abrate and Bonchi (2021); Chhablani et al. (2024); Wellawatte et al. (2022); Verma et al. (2024); Lu et al. (2024). However, this localized perspective constrains their ability to uncover global patterns leveraged by GNNs across multiple graphs, as well as how these patterns integrate into a unified decision-making process. Research on global GNN explainers remains relatively limited Yuan et al. (2020); Huang et al. (2023); Xuanyuan et al. (2023); Azzolin et al. (2023); Armgaan et al. (2024); Fournier and Medya (2025). For instance, XGNN Yuan et al. (2020) and GNNInterpreter Wang and Shen (2023) adopt generative modeling approaches, producing graphs that most strongly correspond to a given class label. For more details, please refer to this survey on GNN explainers Kakkad et al. (2023b).
Table 4: Statistics of the graph datasets.
@@ Appendix B Datasets
B.1 Datasets Details
  1. 1.
    Cora: It is a citation network, in which nodes represent computer science research papers, and each edge between two nodes represents a research paper citing another. The nodes are classified into one of seven categories. Though a citation network is a directed network, the dataset is widely used as an undirected network in the message-passing based graph machine learning, especially for the node classification task, in which the task is to predict each node’s category based on its own text features and text features of its neighbors. Each node’s feature is a 2000-dimensional bag-of-words representation of the keywords in the paper that it corresponds to.
  2. 2.
    WikiCS: It is a text-attributed graph dataset, derived from the Wikipedia platform, widely used for node classification tasks. The nodes correspond to Wikipedia page descriptions of different computer science topics, and edges between nodes represent hyperlinks from one article to another. Each node in the dataset belongs to one of 10 categories. The node classification task is to correctly predict a node’s label, and the edges are considered to be undirected. Each node’s feature vector is a GloVe embedding of size 300, computed based on the text associated with the nodes.
  3. 3.
    Liar: It is a fake-news detection dataset that is often represented as a knowledge graph, with nodes corresponding to statements, speakers, and topics, and edges encoding typed relations such as spoken_by and about.To adapt Liar   into a homogeneous graph suitable for standard GNN pipelines, we merge the three node types—statements, speakers, and topics—into a single unified node set. Each node is embedded with a 384-dimensional BERT representation and augmented with a 3-dimensional one-hot vector indicating its type. Edges representing the original heterogeneous relations (spoken_by and about) are converted to undirected edges and unified into a single edge type in the homogeneous graph. Statement nodes retain their ground-truth labels for fake news classification (ranging from pants-fire to true), while speaker and topic nodes are assigned a dummy label. We also preserve a node-level string attribute—containing the raw statement, speaker name, or topic—for use in explanation and visualization.
  4. 4.
    Amazon-Product: It is a network with nodes representing different kinds of products on Amazon and edges connecting co-purchased products. The nodes belong to one of 47 different categories of products. We use only a subset of the Amazon-Product   dataset, consisting of the first 1000 products and their co-purchase edges, as the entire dataset is very large, and call it Amazon   in our work. The node classification task is to predict a product’s category based on its own as well as its neighbours’ features.
B.2 Relevance of datasets
Table 5: Dataset Descriptions and Relevance to Our Work
@@ B.2 Relevance of datasets
The datasets used in our work are selected based on their relevance to graph explainability tasks, allowing us to assess the effectiveness and robustness of our methods across a diverse set of graph structures and features. The chosen datasets span a variety of domains, including citation networks, text-attributed graphs, fake-news detection, and e-commerce, which reflect both real-world complexity and common challenges faced in explainability. The characteristics of each dataset make them suitable for evaluating the performance of our explanation method. Table 5 summarizes the key features of the datasets and explains their relevance.
Appendix C Additional Implementation Details
We discuss additional details related to the training of the base GNN.
C.1 GNN training

### v3

@@ 5.3 Qualitative Evaluation
Expert Assessment.
We asked experts to evaluate the explanations generated by GNNExplainer and Gspell on the Amazon-Product and Cora datasets over 10 node classification tasks. We follow the qualitative evaluation framework provided in Cedro and Martens (2025), and refer to Appendix E for a full description of the evaluation dimensions. The evaluation framework and setup are presented in Appendix F.
$k=1$ $k=5$ $k=10$ $k=20$ $k=50$
Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size
Amazon Gspell 92.0 4.22 91.1 4.25 88.0 1.71 85.0 1.17 82.5 1.00
Gspell+ 90.5 4.45 92.0 4.24 86.0 1.45 85.0 1.13 81.0 1.01
Liar Gspell 100.0 1.13 100.0 1.11 100.0 1.11 100.0 1.00 OOM OOM
Gspell+ 100.0 1.14 100.0 1.10 98.4 1.10 100.0 1.00 OOM OOM
Cora Gspell 86.2 1.98 86.5 2.07 87.8 2.04 85.4 1.81 84.4 1.33
Gspell+ 85.2 1.87 85.8 2.10 88.0 2.17 87.4 2.17 84.8 1.30
Table 4: Effect of the number of tokens ($k$) on the performance of Gspell and Gspell+.
As Table 2 shows, Gspell’s explanations on Cora were more trustworthy and convincing, likely due to the LLM component’s semantic understanding. However, these explanations were seen as less insightful, suggesting more limited exploration beyond the article’s main subject. On Amazon-Product, Gspell provided more insightful explanations than GNNExplainer by capturing richer semantic links across product categories, but it was rated lower in trustworthiness and convincingness, indicating less consistency and reliability.
Overall, our method performs better on Cora, where its explanations were viewed as more trustworthy, convincing, and communicable. On the more challenging Amazon-Product dataset, GNNExplainer retained an edge in clarity and user confidence, suggesting that dataset complexity and semantic richness shape the relative strengths of LLM-based explanations.
Real Example.
@@
Figure 2 illustrates a prompting example of Gspell on the Amazon dataset to emphasize the narrative quality and interpretability of the generated explanation. In this instance, the target node corresponds to a product (ID 160), predicted to belong to the “Sports & Outdoors” category. Gspell leverages latent GNN representations via soft prompts to benefit from the LLM’s language-generation capabilities while accounting for the local neighborhood of the target node.
Gspell generates node-level rationales from the target node’s computation tree, providing for each neighbor a binary support judgment (YES/NO) indicating whether it supports the target’s classification. Also, it provides a general summary that highlights common patterns across the target’s context, helping to explain the model’s prediction.
6 Ablation Study and Runtime Analysis
6.1 Effect of the projection tokens
Without accounting for GNN internals (embeddings), asking an LLM to generate explanations only based on the input and the predictions of the GNN is theoretically wrong. The main reason is that one can directly substitute the predictive GNN with any arbitrary function that makes predictions based on different reasoning (logic) than the GNN, while still making the same predictions. Therefore, we use the soft prompt generated by the projector for the faithfulness of the explanations.
Soft prompts channel information from the GNN to the LLM, with the token count, $k$, being a critical factor here. Insufficient tokens constrain the information channel, causing the LLM to ignore GNN-derived context in favor of textual priors, thereby increasing hallucination rates. Conversely, a large number of tokens can disrupt coherence during generation. If the soft prompt distribution aligns poorly with the LLM’s embedding space—often due to a suboptimal projection function–increasing $k$ injects out-of-distribution noise into the model. This results in degraded generation quality and increases computational overhead.
We perform an ablation study over different token counts ($k$), shown in Table 4. We observe that $k$ acts as a sparsity regulator. On the complex Amazon dataset, low $k$ yields high fidelity with larger explanations, while high $k$ significantly reduces size at the cost of fidelity. This supports our hypothesis that excessive soft tokens inject noise into the LLM. We find an optimal balance at $k=5$, achieving high fidelity with reduced explanation size. In contrast, Liar, a more heterogeneous dataset, benefits from a higher $k$ as it likely requires more tokens to encode complex relationships than standard homogeneous graphs. Conversely, Cora exhibits high robustness in both fidelity and sparsity across all $k$ values, likely due to its higher homophily.
6.2 Runtime
We provide the details of running times in Table 5, comparing our method with all baselines. Although Gspell is more time-consuming than other baselines, we note that this is justified by the fact that Gspell, unlike the fast baselines, e.g., PGExplainer and TAGExplainer, generates descriptive, informative explanations in natural language that are understandable by humans.
Amazon Cora Liar WikiCS
GNNExplainer 1.45 1.69 1.79 3.76
PGExplainer 0.02 0.03 0.79 0.80
TAGExplainer 0.05 0.13 0.52 1.57
LLM 33.74 18.21 12.80 61.67
Gspell (Ours) 39.65 30.77 42.41 76.76
Table 5: Inference runtime per node for explanation generation measured in seconds.
@@ 7 Conclusions
In this work, we propose Gspell, a lightweight, post-hoc explanation framework that uses LLMs to generate faithful and human-interpretable explanations for GNN predictions on text-attributed graphs. By projecting GNN embeddings into the LLM space, Gspell enables context-aware, natural language rationales without relying on an external GNN explainer. Our approach achieves a superior trade-off between conciseness and fidelity, as well as greater semantic clarity, compared to prior methods. Gspell demonstrates that LLMs can serve as effective interpreters for GNNs and offers a scalable and intuitive path toward explainability in graph-based applications.
Future Work. Although Gspell+ incorporates an InfoNCE loss to enhance faithfulness, it does not consistently outperform Gspell. Moreover, the other projector objectives focus on aligning GNN and LLM embedding spaces rather than directly optimizing explanation quality. Designing more principled, explanation-centered objectives is an important direction for future work.
Additionally, as the computation tree grows, the LLM struggles to capture the full graph context. While batch-wise prompting mitigates this issue, it limits global reasoning. Developing prompt compression techniques that preserve structural information could improve the quality of the generated explanations.
Acknowledgment
The authors acknowledge the National Artificial Intelligence Research Resource (NAIRR) Pilot and the Texas Advanced Computing Center (TACC) Vista for contributing to this research.
@@ 8 Ethical considerations
Potential Risks. While Gspell aims to improve the interpretability of GNN predictions, the generated explanations rely on LLM outputs and may not always faithfully reflect the underlying model’s true decision process. This could lead to over-trust in the explanations. Also, the use of LLMs introduces the risk of hallucinated or misleading reasoning, despite our mitigation strategies. We encourage careful human oversight when deploying such explanation systems in practice.
@@ 9 Limitations
  • Our method combines projected GNN embeddings with text in hybrid prompts. When the computation tree is large, processing the full context can lead to hallucinations, as the LLM struggles to distinguish among multiple soft prompt representations, while batch-wise prompting limits its ability to capture a global view.
  • The explanations generated by Gspell are produced by an LLM conditioned on projected embeddings rather than explicitly tracing the GNN’s computation. Consequently, they may not reflect the true causal factors underlying the model’s prediction, so we do not provide formal guarantees of causal faithfulness.
  • Although the projector aligns GNN embeddings with the LLM embedding space, its training objectives are not directly tailored for explanation. While we explore an InfoNCE-based objective to improve faithfulness, it does not consistently yield gains, suggesting that more effective explanation-specific objectives are needed.
@@ Appendix A Additional Related Work
Here we provide additional related works that are on the integrated methods of LLMs and GNNs. The integration of LLMs with GNNs aims to leverage the advanced reasoning abilities of LLMs and the extensive background knowledge they acquire, alongside the GNNs’ proven capacity to enhance the exploitation of graph structures via message-passing mechanisms Kipf and Welling (2017a); Hamilton et al. (2017). The approaches for integrating LLMs and GNNs primarily vary in the sequence of component application Ren et al. (2024); Chen et al. (2024b).
@@
LLMs as prefix and postfix. First, there are methods where LLMs are used in the prefix modules for task and feature representation Liu et al. (2024a); Fang et al. (2024); He et al. (2024a); Yang et al. (2025). This architecture provides greater flexibility, as various tasks can be specified in natural language with text-based graph representations, and LLM modules can remain fixed. On the other hand, the second class of methods uses GNNs as the prefix Wang et al. (2023); Chen et al. (2024a); Hu et al. (2024); Tang et al. (2024), which is better suited for capturing the graph topology and its intricate relationships and uses LLMs for prediction. This approach also has some flexibility, as the LLM can output embeddings, prediction scores, or textual sequences to address various downstream problems specifically for generation tasks Wang et al. (2023); Chen et al. (2024a); Hu et al. (2024); Tang et al. (2024).
@@
Independent LLMs and joint LLM-GNNs. Additionally, there is a body of work that employs LLMs independently to tackle graph-related tasks Cheng et al. (2024); Fatemi et al. (2024); Li et al. (2024). Approaches following this paradigm benefit from LLMs’ generalization capabilities at both the input and output levels and can potentially avoid training learnable parameters by fixing the LLM component and relying on input prompts. Finally, there is a group of works which employ Gnns and LLMs jointly through soft prompting, which necessitates alignments between their embeddings while enhancing LLMs reasoning by the Gnn captured graph structural features Liu et al. (2024b); Wang et al. (2024); Tian et al. (2023). However, these works do not address the GNN explanation problem.
@@
Explainers for GNNs. Several algorithms have been proposed in the literature to enhance the interpretability of GNNs. The majority of existing explainers concentrate on instance-level explanations. Instance-level (or local) explainers operate on individual graphs, identifying influential components—such as subgraphs—that most strongly impact the model’s prediction Ying et al. (2019b); Luo et al. (2020); Shan et al. (2021); Huang et al. (2022); Yuan et al. (2022); Lucic et al. (2022b); Tan et al. (2022b); Lin et al. (2021); Bajaj et al. (2021); Abrate and Bonchi (2021); Chhablani et al. (2024); Wellawatte et al. (2022); Verma et al. (2024); Lu et al. (2024). However, this localized perspective constrains their ability to uncover global patterns leveraged by GNNs across multiple graphs, as well as how these patterns integrate into a unified decision-making process. Research on global GNN explainers remains relatively limited Yuan et al. (2020); Huang et al. (2023); Xuanyuan et al. (2023); Azzolin et al. (2023); Armgaan et al. (2024); Fournier and Medya (2025). For instance, XGNN Yuan et al. (2020) and GNNInterpreter Wang and Shen (2023) adopt generative modeling approaches, producing graphs that most strongly correspond to a given class label. For more details, please refer to this survey on GNN explainers Kakkad et al. (2023).
@@ Appendix B Datasets
Table 7: Dataset Descriptions and Relevance to Our Work
B.1 Datasets Details
  1. 1.
    Cora: It is a citation network, in which nodes represent computer science research papers, and each edge between two nodes represents a research paper citing another. The nodes are classified into one of seven categories. Though a citation network is a directed network, the dataset is widely used as an undirected network in the message-passing-based graph machine learning, especially for the node classification task, in which the task is to predict each node’s category based on its own text features and text features of its neighbors. Each node’s feature is a 2000-dimensional bag-of-words representation of the keywords in the paper that it corresponds to.
  2. 2.
    WikiCS: It is a text-attributed graph dataset, derived from the Wikipedia platform, widely used for node classification tasks. The nodes correspond to Wikipedia page descriptions of different computer science topics, and the edges represent hyperlinks from one article to another. Each node in the dataset belongs to one of 10 categories. The node classification task is to correctly predict a node’s label, and edges are undirected. Each node’s feature vector is a 300-dimensional GloVe embedding computed from the text associated with the node.
  3. 3.
    Liar: It is a fake-news detection dataset that is often represented as a knowledge graph, with nodes corresponding to statements, speakers, and topics, and edges encoding typed relations such as spoken_by and about. To adapt Liar   into a homogeneous graph suitable for standard GNN pipelines, we merge the three node types—statements, speakers, and topics—into a single unified node set. Each node is embedded with a 384-dimensional BERT representation and augmented with a 3-dimensional one-hot vector indicating its type. Edges representing the original heterogeneous relations (spoken_by and about) are converted to undirected edges and unified into a single edge type in the homogeneous graph. Statement nodes retain their ground-truth labels for fake news classification (ranging from pants-fire to true), while speaker and topic nodes are assigned a dummy label. We also preserve a node-level string attribute containing the raw statement, speaker name, or topic for use in explanations and visualizations.
  4. 4.
    Amazon-Product: It is a network with nodes representing different product categories on Amazon and edges connecting co-purchased products. The nodes belong to one of 47 product categories. We use only a subset of the Amazon-Product   dataset, consisting of the first 1000 products and their co-purchase edges, as the entire dataset is very large, and call it Amazon   in our work. The node classification task is to predict a product’s category based on its own features and those of its neighbors. For nodes, bag-of-words representations are derived from node textual attributes.
B.2 Relevance of datasets
@@
The datasets used in our work are selected for their relevance to graph explainability tasks, enabling us to assess the effectiveness and robustness of our methods across a diverse set of graph structures and features. The chosen datasets span a variety of domains, including citation networks, text-attributed graphs, fake news detection, and e-commerce, and reflect both real-world complexity and common challenges in explainability. The characteristics of each dataset make them suitable for evaluating the performance of our explanation method. Table 7 summarizes the key features of the datasets and explains their relevance.
Appendix C Extended Experiments
C.1 GNN Architecture
We extend Table 1 of the paper with additional experiments for Amazon using two other standard GNN architectures: GIN Xu et al. and GAT Veličković et al. (2017). The results are reported in Table 8.
Method Model Fidelity (%) Size
GNNExplainer GAT 100.0 10.57
PGExplainer 71.8 4.45
TAGExplainer 63.1 1.03
Random 89.9 7.28
LLM 96.5 4.92
Gspell 86.8 3.79
GNNExplainer GIN 100.0 10.57
PGExplainer 87.2 5.15
TAGExplainer 30.9 1.03
Random 84.7 7.28
LLM 86.0 4.90
Gspell 73.7 3.06
Table 8: Performance comparison on the Products dataset using different GAT and GIN as two other architectures (using "Llama 3.1 8B Instruct"). Higher fidelity and lower size are better.
With GCN architecture, we observed that our method produces the most compact explanations (close to size 1) across all architectures, achieving a strong balance. However, while explanation size remains minimal for the GAT and GIN models, fidelity drops more noticeably, particularly with GIN (73.7%), suggesting that our projection-based mechanism may be less effective when the model relies heavily on structural information or deep aggregation. In contrast, GNNExplainer maintains the highest fidelity across all architectures, but at the cost of much larger explanations, while PGExplainer and TAGExplainer underperform in fidelity despite moderate sizes. The LLM baseline shows robust fidelity, especially on GAT, but its explanations are substantially less concise. These results suggest that our method offers an excellent fidelity/interpretability trade-off on GCN, highlighting the need for architecture-aware extensions when applied to more structure-based models.
Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size Fidelity (%) Size
LLM 86.0 4.90 88.0 5.14 84.0 2.33 88.0 5.17 84.0 2.34
Gspell 91.1 4.25 83.0 1.63 80.0 1.32 80.0 1.47 80.0 1.12
C.2 LLM Backbone
We also evaluate our method across different LLM models and present the results in Table 9. The baseline LLM method, which directly leverages the language model without projecting GNN embeddings, generally achieves higher fidelity scores but produces substantially larger explanations, two to three times larger. In contrast, our method consistently generates more compact explanations—typically around size 1 —while maintaining comparable fidelity. Notably, Llama 3.1 and Mistral backbones show strong fidelity across both methods, whereas smaller models such as GPT-Neo, Phi-3 Mini, and Pythia exhibit lower fidelity overall. These results highlight our method’s effectiveness at generating precise explanations while maintaining reasonable alignment with the GNN’s predictions and demonstrate its compatibility with different LLM models.
C.3 Framework Component Analysis
To assess the effectiveness of the different components of our approach, we measure the performance of our method with and without the projector and the post-processing. The results are in Table 10. We observe that both fidelity and explanation size decrease across all datasets when we replace the prompt words with the projected GNN embeddings. This might be due to the LLM having to extract knowledge from more abstract representations, making it more selective in its reasoning. We also observe that adding the post-processing step increases fidelity at a lower cost in size than the method without a projector on the WikiCS dataset. From this analysis, we suggest that our method offers a good compromise: achieving high fidelity (slightly lower than directly prompting the LLM with node-associated text) while maintaining a compact explanation size (slightly larger than projected GNN embeddings without post-processing).
Amazon Cora Liar WikiCS
Fidelity Size Fidelity Size Fidelity Size Fidelity Size
LLM 86.0 4.90 94.5 3.27 100.0 2.22 91.4 21.14
LLM+Pr 85.8 3.2 94.5 3.27 100.0 1.10 80.2 3.7
LLM+Pr+Po 91.1 4.25 86.5 2.07 100.0 1.11 86.5 12.8
Appendix D Additional Implementation Details
We provide additional details of the experiments as follows:
D.1 Experimental Setup
Datasets.
For all datasets, we use the graph topology as provided, without any modification or filtering of nodes or classes. For AMAZON-PRODUCT, we first randomly subsample 1000 nodes from the full dataset. For all datasets, we use a consistent split protocol with 60% of nodes for training, 10% for validation, and 200 held-out nodes for testing. Across all datasets, we restrict computation trees to at most 2 hops from each node.
Models.